OpenStudy (anonymous):

I need help simplifying the expression radical3sqrt16e^12

jhonyy9 (jhonyy9):

3sqrt(16e^12) = 3*(16e^12)^(1/2) = 3*(4^2 *(e^6)^2)^(1/2) = 3*((4*(e^6))^2^(1/2) = = 3*(4*e^6)^(2*(1/2)) = 3*(4*e^6)^(?) = ?